You are here

Observation

Status: 
Wykonane
Observation time: 
Monday, April 30, 2012 - 21:00 to Tuesday, May 1, 2012 - 10:30
Observer: 
Instruktor: 
Objects: 
Spectra of: SAO 81004,α Oph,β Oph,β Lyr,ε Lyr,Vega, SAO 106316
Is scope parked?: 
Tak
Is the dome parked?: 
Tak

The spectrograph is mounted on the telescope.

The photometric filter is placed inside the CCD.

Images: